# From Local to Global Order: A Theory of Neural Synaptic Balance

### v2

@@ Abstract.
We develop a general theory of neural synaptic balance and how it can emerge or be enforced in neural networks. For a given additive cost function $R$ (regularizer), a neuron is said to be in balance if the total cost of its input weights is equal to the total cost of its output weights. The basic example is provided by feedforward networks of ReLU units trained with $L_{2}$ regularizers, which exhibit balance after proper training. The theory explains this phenomenon and extends it in several directions. The first direction is the extension to bilinear and other activation functions. The second direction is the extension to more general regularizers, including all $L_{p}$ ($p>0$) regularizers. The third direction is the extension to non-layered architectures, recurrent architectures, convolutional architectures, as well as architectures with mixed activation functions. Gradient descent on the error function alone does not converge in general to a balanced state where every neuron is in balance, even when starting from a balanced state. However, gradient descent on the regularized error function must converge to a balanced state, and thus network balance can be used to assess learning progress. The theory is based on two local neuronal operations: scaling which is commutative, and balancing which is not commutative. Finally, and most importantly, given any initial set of weights, when local balancing operations are applied to each neuron in a stochastic manner, global order always emerges through the convergence of the stochastic balancing algorithm to the same unique set of balanced weights. The reason for this convergence is the existence of an underlying strictly convex optimization problem where the relevant variables are constrained to a linear, only architecture-dependent, manifold. The theory is corroborated through various simulations carried out on benchmark data sets. Scaling and balancing operations are entirely local and thus physically plausible in biological and neuromorphic networks.
@@ 1. Introduction
One of the most common complaints against neural networks is that they provide at best “black-box” solutions to problems. In particular, when large neural networks are trained on complex tasks, they produce large arrays of synaptic weights that have no clear structure and are difficult to interpret. Thus finding any kind of structure in the weights of large neural networks is of great interest. Here we study a particular kind of structure we call neural synaptic balance and the conditions under which it emerges. Neural synaptic balance is different from the biological notion of balance between excitation and inhibition [12, 10, 13, 17, 23]. We use the term to refer to any systematic relationship between the input and output synaptic weights of individual neurons or layers of neurons. Here we consider the case where the cost of the input weights is equal to the cost of the output weights, where the cost is defined by some regularizer. One of the most basic examples of such a relationship is when the sum of the squares of the input weights of a neuron is equal to the sum of the squares of its output weights.
@@
where $IN$ and $OUT$ denote the set of incoming and outgoing weights respectively, $A=\sum_{i\in IN}w_{i}^{2}$, and $B=\sum_{i\in OUT}w_{i}^{2}$. The product of the two terms on the right-hand side of Equation 1.1 is equal to $AB$ and does not depend on $\lambda$. Thus, the minimum is achieved when these two terms are equal, which yields: $(\lambda^{*})^{4}=B/A$ for the optimal $\lambda^{*}$. The corresponding new set of weights, $v_{i}=\lambda^{*}w_{i}$ for the input weights and $v_{i}=w_{i}/\lambda^{*}$ for the outgoing weights, must be balanced: $\sum_{i\in IN}v_{i}^{2}=\sum_{i\in OUT}v_{i}^{2}$. This is because the optimal scaling factor for these rescaled weights must be $\lambda^{*}=1$. Furthermore, if an entire network of ReLU neurons is properly trained using a standard error function with an $L_{2}$ regularizer, at the end of training one observes a remarkable phenomenon: for each ReLU neuron, the norm of the incoming synaptic weights is approximately equal to the norm of the outgoing synaptic weights, i.e. every neuron is balanced.
@@
There have been isolated previous studies of this kind of synaptic balance [9, 25] under special conditions. For instance, in [9], it is shown that if a deep network is initialized in a balanced state with respect to the sum of squares metric, and if training progresses with an infinitesimal learning rate, then balance is preserved throughout training. Here, we take a different approach aimed at uncovering the generality of neuronal balance phenomena, the learning conditions under which they occur, as well as new local balancing algorithms and their convergence properties. We explain and study neural synaptic balance in its generality in terms of activation functions, regularizers, network architectures, and training stages. In particular, we systematically answer questions such as: Why does balance occur? Does it occur only with ReLU neurons? Does it occur only with $L_{2}$ regularizers? Does it occur only in fully connected feedforward architectures (as opposed to, for instance, locally connected, convolutional, or recurrent architectures)? Does it occur only at the end of training? In the process of answering these questions, we introduce local scaling and balancing operations for individual neurons or entire neural layers. Furthermore, we show that when these local operations are applied stochastically, a global balanced state always emerges, and this state is unique and depends only on the initial weights, but not on the order in which the neurons are balanced.
@@ 2.1. Additive Activation Functions
A neuronal activation function $f:\mathbb{R}\to\mathbb{R}$ is additively linear if and only if $f(x+y)=f(x)=(f(y)$ for any real numbers $x$ and $y$.
@@ 5. General Framework and Single Neuron Balance
Under the assumptions of the theorem, $E$ is unchanged under the rescaling of the incoming and outgoing weights of unit $i$ due to the homogeneity property of BiLUs. Without any loss of generality, let us assume that at the beginning: $\sum_{w\in IN(i)}g_{w}(w)<\sum_{w\in OUT(i)}g_{w}(w)$. As we increase $\lambda$ from 1 to +$\infty$+ ∞, by the assumptions on the functions $g_{w}$, the term $\sum_{w\in IN(i)}g_{w}(\lambda w)$ increases continuously from its initial value to +$\infty$+ ∞, whereas the term $\sum_{w\in OUT(i)}g_{w})w/\lambda)$ decreases continuously from its initial value to 00. Thus, there is a unique value $\lambda^{*}$ where the balance is realized. If at the beginning $\sum_{w\in IN(i)}g_{w}(w)>\sum_{w\in OUT(i)}g_{w}(w)$, then the same argument is applied by decreasing $\lambda$ from 1 to 0. ∎
@@ 6.1. Bi-Power Units (BiPU)
Given the general shape, these activations functions can be called BiPU (Bi-Power-Units). Note that in the general case where $c>0$, $C$ and $D$ do not need to be equal. In particular, one of them can be equal to zero, and the other one can be different from zero giving rise to “rectified power units” (Figure 1).
Figure 1.
@@ 9.2. Convergence of the Basic Stochastic Balancing Algorithm to a Unique Optimum
Thus the constraints associated with being a self-consistent configuration of $L_{ij}$’ s are all linear. This linear manifold of constraints depends only on the architecture, i.e., the graph of connections. The strictly convex function $R(L_{ij})$ depends on the actual weights $W$. Different sets of weights $W$ produce different convex functions over the same linear manifold.
@@
The same convergence to the unique global optimum is observed if each neuron, when stochastically visited, is favorably scaled rather than balanced, i.e., it is scaled with a factor that reduces $R$ but not necessarily minimizes $R$. Stochastic balancing can also be viewed as a form of EM algorithm where the E and M steps can be taken fully or partially.
@@
Figure 7. SGD applied to $E$ alone, in general, does not converge to a balanced state, but sGD applied to $E+R$ converges to a balanced state. (A-C) Simulations use a deep fully connected autoencoder trained on the MNIST dataset. (D-F) Simulations use a deep locally connected network trained on the CFAR10 dataset. (A,D) Regularization leads to neural balance. (B,E) The training loss decreases and converges during training (these panels are not meant for assessing the quality of learning when using a regularizer). (C,F) Using weight regularization decreases the norm of weights. (A-F) Shaded areas correspond to one s.t.d around the mean (in some cases the s.t.d. is small and the shaded area is not visible).
@@
Figure 8. Even if the starting state is balanced, SGD does not preserve the balance unless the learning rate is infinitely small. (A-C) Simulations use a deep fully connected autoencoder trained on the MNIST dataset. (D-F) Simulations use deep locally connected network trained on the CFAR10 dataset. (A-F) The initial weights are balanced using the stochastic balancing algorithm. Then the network is trained by SGD. (A,D) When the learning rate (lr) is relatively large, without regularization, the initial balance of the network is rapidly disrupted. (B,E) The training loss decreases and converges during training (these panels are not meant for assessing the quality of learning when using a regularizer). (C,F) Using weight regularization decreases the norm of the weights. (A-F) Shaded areas correspond to one s.t.d around the mean (in some cases the s.t.d. is small and the shaded area is not visible).
10. Simulations
@@
To further corroborate the results, we ran multiple experiments. Here we report the results from two series of experiments. The first one is conducted using a six-layer, fully connected, autoencoder trained on MNIST [8] for a reconstruction task with ReLU activation functions in all layers and the sum of squares errors loss function. The number of neurons in consecutive layers, from input to output, is 784, 200, 100, 50, 100, 200, 784. Stochastic gradient descent (SGD) learning by backpropagation is used for learning with a batch size of 200.
@@
The second one is conducted using three locally connected layers followed by three fully connected layers trained on CFAR10 [18] for a classification task with leaky ReLU activation functions in the hidden layers, a softmax output layer, and the cross entropy loss function. The number of neurons in consecutive layers, from input to output, is 3072, 5000, 2592, 1296, 300, 100, 10. Stochastic gradient descent (SGD) learning by backpropagation is used for learning with a batch size of 5.
@@
Code Availability: The code for reproducing the simulation results is available under the Apache 2.0 license at: https://github.com/ARahmansetayesh/a-theory-of-neural-synaptic-balance.

### v3

# A Theory of Synaptic Neural Balance: From Local to Global Order

@@ Abstract.
We develop a general theory of synaptic neural balance and how it can emerge or be enforced in neural networks. For a given additive cost function $R$ (regularizer), a neuron is said to be in balance if the total cost of its input weights is equal to the total cost of its output weights. The basic example is provided by feedforward networks of ReLU units trained with $L_{2}$ regularizers, which exhibit balance after proper training. The theory explains this phenomenon and extends it in several directions. The first direction is the extension to bilinear and other activation functions. The second direction is the extension to more general regularizers, including all $L_{p}$ ($p>0$) regularizers. The third direction is the extension to non-layered architectures, recurrent architectures, convolutional architectures, as well as architectures with mixed activation functions. Gradient descent on the error function alone does not converge in general to a balanced state, where every neuron is in balance, even when starting from a balanced state. However, gradient descent on the regularized error function ought to converge to a balanced state, and thus network balance can be used to assess learning progress. The theory is based on two local neuronal operations: scaling which is commutative, and balancing which is not commutative. Finally, and most importantly, given any initial set of weights, when local balancing operations are applied to each neuron in a stochastic manner, global order always emerges through the convergence of the stochastic balancing algorithm to the same unique set of balanced weights. The reason for this convergence is the existence of an underlying strictly convex optimization problem where the relevant variables are constrained to a linear, only architecture-dependent, manifold. The theory is corroborated through various simulations carried out on benchmark data sets. Additional simulations show that balancing neurons prior to learning, or during learning in alternation with gradient descent steps, can improve learning speed and final performance thereby expanding the arsenal of available training tools. Scaling and balancing operations are entirely local and thus physically plausible in biological and neuromorphic networks.
@@ 1. Introduction
One of the most common complaints against neural networks is that they provide at best “black-box” solutions to problems. In particular, when large neural networks are trained on complex tasks, they produce large arrays of synaptic weights that have no clear structure and are difficult to interpret. Thus finding any kind of structure in the weights of large neural networks is of great interest. Here we study a particular kind of structure we call synaptic neural balance and the conditions under which it emerges. Neural balance in general refers to some kind of homeostatic process that facilitates information processing in neural systems. Synaptic neural balance is different from other forms of neural balance which are based on the connectivity patterns or the activities of neurons, such as the biological notion of balance between excitation and inhibition [14, 12, 15, 19, 28]. In contrast, we use the term synpatic neural balance to refer to any systematic relationship between the input and output synaptic weights of individual neurons, or layers of neurons. Here we consider the case where the cost of the input weights is equal to the cost of the output weights, where the cost is defined by some cost function or regularizer. One of the most basic examples of such a relationship is when the sum of the squares of the input weights of a neuron is equal to the sum of the squares of its output weights.
@@
where $IN$ and $OUT$ denote the set of incoming and outgoing weights respectively, $A=\sum_{i\in IN}w_{i}^{2}$, and $B=\sum_{i\in OUT}w_{i}^{2}$. The product of the two terms on the right-hand side of Equation 1.1 is equal to $AB$ and does not depend on $\lambda$. Since of all the rectangles with the same area the square has the shortest perimeter, the minimum is achieved when these two terms are equal, which yields: $(\lambda^{*})^{4}=B/A$ for the optimal $\lambda^{*}$. The corresponding new set of weights, $v_{i}=\lambda^{*}w_{i}$ for the input weights and $v_{i}=w_{i}/\lambda^{*}$ for the outgoing weights, must be balanced: $\sum_{i\in IN}v_{i}^{2}=\sum_{i\in OUT}v_{i}^{2}$. This is because the optimal scaling factor for these rescaled weights must be $\lambda^{*}=1$. Furthermore, if an entire network of ReLU neurons is properly trained using a standard error function with an $L_{2}$ regularizer, at the end of training one observes a remarkable phenomenon: for each ReLU neuron, the norm of the incoming synaptic weights is approximately equal to the norm of the outgoing synaptic weights, i.e. every neuron is balanced. Our goal here is to gain a deeper understanding of this basic example and of how general the phenomenon of synaptic balance is.
There have been isolated previous studies of this kind of balance [11, 30] from different perspectives or under special conditions, restricted to particular neuronal models, architectures, or cost functions. Many of these studies propose to favor balance by adding an extra term to the loss function forcing the difference between input and output synaptic costs to be close to zero, or the ratio between input and output synaptic costs to be close to one [36]. The work in [30] proposes a new biologically plausible learning rule to increase robustness of neural dynamics in non-linear recurrent networks resulting in a form of synaptic balance. In [11], the authors show that if a deep network is initialized in a balanced state with respect to the sum of squares metric, and if training progresses with an infinitesimal learning rate, then the balance is preserved throughout training. Related results are also described in [2]. Other studies have also explored symmetry and balance effects on training neural networks. For example, [21] shows that training with stochastic gradient descent does not work well in highly unbalanced neural networks. As a result, these authors propose a rescaling-invariant solution analyzed in [23]. In the same vein, other authors have proposed that learning in neural networks can be accelerated with rescaling transformations [37, 1] without focusing on synaptic balancing. In [27] multiplicative rescaling factors, one at each hidden unit, are used to balance the weights with a proof of convergence, but not of uniqueness.
@@
Here we take a different, more unified, approach aimed at uncovering the generality of synaptic neuronal balance phenomena, the learning conditions under which they occur, as well as new local balancing algorithms and their convergence properties. We explain and study synaptic neural balance in its generality in terms of activation functions, regularizers, network architectures, and training stages. In particular, we systematically answer questions such as: Why does balance occur? Does it occur only with ReLU neurons? Does it occur only with $L_{2}$ regularizers? Does it occur only in certain architectures, e.g., fully connected feedforward architectures, as opposed to locally connected, convolutional, or recurrent architectures? Does it occur only at the end of training? In the process of answering these questions, we introduce local scaling and balancing operations for individual neurons or entire neural layers. Furthermore, we show that when these local operations are applied stochastically, a global balanced state always emerges, and this state is unique and depends only on the initial weights, but not on the order in which the neurons are balanced. In one rare case, the state can be solved analytically (Appendix). In addition, we provide experimental results showing that balancing neurons, before or during training, can improve learning convergence and performance in both feedforward and recurrent networks.
@@ 2.1. Additive Activation Functions
A neuronal activation function $f:\mathbb{R}\to\mathbb{R}$ is additively linear if and only if $f(x+y)=f(x)+(f(y)$ for any real numbers $x$ and $y$.
@@ 5. General Framework and Single Neuron Balance
Under the assumptions of the theorem, $E$ is unchanged under the rescaling of the incoming and outgoing weights of unit $i$ due to the homogeneity property of BiLUs. Without any loss of generality, we assume that at the beginning: $\sum_{w\in IN(i)}g_{w}(w)<\sum_{w\in OUT(i)}g_{w}(w)$. As we increase $\lambda$ from 1 to +$\infty$+ ∞, by assumptions on functions $g_{w}$, the term $\sum_{w\in IN(i)}g_{w}(\lambda w)$ increases continuously from its initial value to +$\infty$+ ∞, whereas the term $\sum_{w\in OUT(i)}g_{w})w/\lambda)$ decreases continuously from its initial value to 00. Thus, there is a unique value $\lambda^{*}$ where the balance is realized. If at the beginning $\sum_{w\in IN(i)}g_{w}(w)>\sum_{w\in OUT(i)}g_{w}(w)$, then the same argument is applied by decreasing $\lambda$ from 1 to 0. ∎
@@ 6.1. Bi-Power Units (BiPU)
Given the general shape, these activation functions can be called BiPU (Bi-Power-Units). Note that in the general case where $c>0$, $C$ and $D$ do not need to be equal. In particular, one of them can be equal to zero, and the other one can be different from zero giving rise to “rectified power units” (Figure 1).
Figure 1. BiPU activation functions (Bi-Power-Units) as described in Equation 6.3
@@ 9.2. Convergence of the Basic Stochastic Balancing Algorithm to a Unique Optimum
Thus the constraints associated with being a self-consistent configuration of $L_{ij}$’ s are all linear. This resulting linear manifold $\mathcal{L}$ depends only on the architecture, i.e., the graph of connections, but not on the actual weight values. The strictly convex function $R(L_{ij})$ depends on the actual weights $W$. Different sets of weights $W$ produce different convex functions over the same linear manifold. If $E$ denotes the total number of connections, then obviously $\dim{\mathcal{L}}\leq E$. In order to infer all the $\Lambda_{i}$, there must exist at least one constrained path going through each node $i$. Thus, in a layered feedforward network, the dimension of $\mathcal{L}$ is given by: $\dim{\mathcal{L}}=E-M$, where here $M$ denotes the size of the largest layer.
@@
Remark 9.7.
All the results in this section remain true in a mixed network containing both BiLU neurons and non-BiLU neurons, as long as one uses $\Lambda_{i}=1$ for any non-BiLU neuron.
@@
The same convergence to the unique global optimum is observed if each neuron, when stochastically visited, is partially balanced (or favorably scaled) rather than fully balanced, i.e., it is scaled with a factor that reduces $R$ but not necessarily minimizes $R$, as long as these factors are not taken to be infinitesimally close to 1. . Stochastic balancing can also be viewed as a form of EM algorithm where the E and M steps can be taken fully or partially. Partial balancing can be faster and is used in some of the experiments of Section 11.
Remark 9.10.
Stochastic balancing of synapses can be applied to any network, even networks comprised entirely of non-homogeneous neurons. In this case, stochastic balancing will still converge to a unique stable configuration of the synaptic weights. However, the overall function implemented by the balanced network may differ from the function implemented by the network at the start of the stochastic balancing. Experiments on balancing sigmoidal neurons are reported in Section 11.
Remark 9.11.
In principle, balancing (or scaling) can be applied independently of the regularization approach being used. For instance, during training, one could alternate scaling and stochastic gradient steps, where the scaling step is performed with respect to $L_{1}$, but the gradient descent step is performed with respect to an $L_{2}$ regularized error function.
@@
Figure 7. SGD applied to $E$ alone, in general, does not converge to a balanced state, but sGD applied to $E+R$ converges to a balanced state. (A-C) Simulations use a deep fully connected autoencoder trained on the MNIST dataset. (D-F) Simulations use a deep locally connected network trained on the CIFAR10 dataset. (A,D) Regularization leads to neural balance. (B,E) The training loss decreases and converges during training (these panels are not meant for assessing the quality of learning when using a regularizer). (C,F) Using weight regularization decreases the norm of weights. (A-F) Shaded areas correspond to one s.t.d around the mean (in some cases the s.t.d. is small and the shaded area is not visible).
@@
Figure 8. Even if the starting state is balanced, SGD does not preserve the balance unless the learning rate is infinitely small. (A-C) Simulations use a deep fully connected autoencoder trained on the MNIST dataset. (D-F) Simulations use deep locally connected network trained on the CIFAR10 dataset. (A-F) The initial weights are balanced using the stochastic balancing algorithm. Then the network is trained by SGD. (A,D) When the learning rate (lr) is relatively large, without regularization, the initial balance of the network is rapidly disrupted. (B,E) The training loss decreases and converges during training (these panels are not meant for assessing the quality of learning when using a regularizer). (C,F) Using weight regularization decreases the norm of the weights. (A-F) Shaded areas correspond to one s.t.d around the mean (in some cases the s.t.d. is small and the shaded area is not visible).
9.3. Convergence to a Unique Optimum for BiPU Stochastic Balancing
We have seen that a generalized form of scaling and balancing can be defined for more general units than BiLUs, in particular for BiPUs. Thus now we consider a network of units with activations functions $f$ satisfying the relationship: $f(\lambda x)=\lambda^{c}f(x)$ (note that this includes BiLU units for $c=1$). We even allow $c$ to vary from unit to unit.
It is easy to see that most of the analyses above done for BiLU units apply to this generalization. In particular, if we apply stocahstic generalized balancing, in the limit the positive multipliers of each connection $w_{ij}$ must satisfy:
$M_{ij}=\Lambda_{i}/\Lambda_{j}^{c_{j}}$ (9.8)
As above, we can define a new set of variables $L_{ij}=\log M_{ij}$ and, for any $p>0$, the regularizer $R(L)=\sum_{ij}e^{pL_{ij}}|w_{ij}|^{p}$ is strictly convex. What is different, however, is the set of constraints on the variables $L_{ij}$. These are the constraints that allow one to compute the variables $\Lambda_{i}$ uniquely from the variables $L_{ij}$ (or, equivalently, the variables $M_{ij}$). This is addressed by the following theorem.
Theorem 9.12.
Under the same conditions of Theorem 9.1, but using activation functions that satisfy for each unit $i$ the relationship $f(\lambda x)=\lambda^{c_{i}}f(x)$, the corresponding stochastic generalized balancing algorithm converges to the unique minimum of a strictly convex optimization problem in the variables $L_{ij}$. The strictly convex objective function is given by $R(L)=\sum_{ij}e^{pL_{ij}}|w_{ij}|^{p}$. The constraints are linear and of the form:
$\sum_{i\in\pi}\left(\prod_{k=i}^{n}c_{k}\right)L_{ii-1}=0$ (9.9)
for each path $\pi$ from an input unit to an output unit, going sequentially through the units $0,1,\ldots,n$, where 0 corresponds to the input unit, and $n$ corresponds to the output unit of the path. The set of paths in the constraints must cover all the units in the network.
Proof.
Let us assume that there is a consistent set of multipliers $\Lambda_{0},\ldots,\Lambda_{n}$ associated with the coefficients $L_{ii-1}=\log M_{ii-1}$ along the path $\pi$, with $\Lambda_{0}=\Lambda_{n}=1$. Since $M_{ii-1}=\Lambda_{i}/\Lambda_{i-1}^{c_{i-1}}$, we can derive the multipliers $\Lambda_{i}$ iteratively by propagating information from the input unit to the output unit, in the form:
$\Lambda_{i}=M_{ii-1}\Lambda_{i-1}^{c_{i-1}}\quad{\rm or}\quad\log\Lambda_{i}=L_{ii-1}+c_{i-1}\log\Lambda_{i-1}$ (9.10)
Using the boundary conditions $\Lambda_{0}=\Lambda_{n}=1$ gives the formula in Theorem 9.12. The same arguments given for BiLU units can be used to complete the proof. ∎
Remark 9.13.
Note that if all the units have the same exponent $c$ associated with the scaling of their activation functions, then the linear constraints have the simplified form:
$\sum_{i\in\pi}c^{n+1-i}L_{ii-1}=0$ (9.11)
Next we present two sets of simulations. The first set of simulations aims to corroborate the theory. The second set of simulations is more practical and shows how balancing can be used in practice during training to improve overall performance.
10. Simulations: Corroboration of the Theory
@@
The code and experiments for this section are publicly available at https://github.com/ARahmansetayesh/a-theory-of-neural-synaptic-balance. To further corroborate the results, we ran multiple experiments. Here we report the results from two series of experiments. The first one is conducted using a six-layer, fully connected, autoencoder trained on MNIST [10] for a reconstruction task with ReLU activation functions in all layers and the sum of squares errors loss function. The number of neurons in consecutive layers, from input to output, is 784, 200, 100, 50, 100, 200, 784. Stochastic gradient descent (SGD) learning by backpropagation is used for learning with a batch size of 200.
@@
The second one is conducted using three locally connected layers followed by three fully connected layers trained on CIFAR10 [20] for a classification task with leaky ReLU activation functions in the hidden layers, a softmax output layer, and the cross entropy loss function. The number of neurons in consecutive layers, from input to output, is 3072, 5000, 2592, 1296, 300, 100, 10. Stochastic gradient descent (SGD) learning by backpropagation is used for learning with a batch size of 5.